UNIVERSITY OF LIVERPOOL Question 4 (a) In Selective Laser Melting (SL.M) steel powder is deposited on a build platform and is melted by a fibre laser. Calculate the energy required to melt Icc of steel powder using the following quantities The density of steel The melting point of stecl Average specific heat capacity of steel The latent heat of fusion for steel Ambient temp Absorbtivity of the laser 7800kg/m 1515°C 620 J/kg K 247 J/g 20°C 0.4 Given below are the process parameters for processing the same steel powder in an actual SLM machine. Calculate the energy to melt 1ce of steel and discuss why the energy needed in an actual SLM is different to that in the Laser Power 200 W Exposure 3.00E-04 s Point dist 7.00E-05 m Scan spacing 1.00E-04 m Layer Thickness 50E-06 m [10 marks] (b) In Selective Laser Sintering (SLS), a polymer powder is processed to produce three dimensional parts. Draw a typical SLS system and describe the environmenta! and powder deposition devices used in the process. You should also indicate the machine features required to make the production of [6 marks] stress free, non-oxidised parts. (e) Name 2 types of materials which can be used in SLS and discuss their [4marks] characteristics.
Answer #1

SLM involves a high power-density laser to melt and fuse metallic powders together.The melting of powder strongly depends on the input of laser energy supplied to the material.The energy required to melt 1cc of steel is given by

E =  Laser power (P)/scanning velocity (v) X scan spacing (s) X Layer thickness (t)

Here,scanning velocity= point distance/exposure

Therefore,E = 200 / ((7x10-5/3x10-4) X (1x10-4) X ( 50x10-6))

= 200 / (0.233) x 5 x 10-9 = 171 x 109 W = 171 x 106 KW